Written Assignment unit 7




Complete the following questions utilizing the concepts introduced in this unit.



1. Find the length of an arc in a circle of radius 10 centimetres subtended by the central angle of 50°. Show your work.





2. Graph f(x) = x Sin x on [-4π, 4π] and verbalize how the graph varies from the graphs of f(x) + or - x.



Graph f(x) =Sin x / x on the window [−5π, 5π] and describe freely what the graph shows. You can use www.desmos.com/calculator to obtain the graphs.




3. A 23-ft ladder leans against a building so that the angle between the ground and the ladder is 80°. How high does the ladder reach up the side of the building? Show the steps of your reasoning.




1
Expert's answer
2022-05-31T13:47:16-0400

1. The circumference of the circle is "L=2\pi r=2\pi(10)=20\pi(cm)"

The length "l" of yhe arc in a circle subtended by the central angle of 50° is


"l=L(\dfrac{50\degree}{360\degree})=\dfrac{20\pi(5)}{36}=\dfrac{25\pi}{9}(cm)"

2.

First draw the graph of "f(x)=\u00b1x."

The function "f(x)=x\sin x" is even. The graph is symmetric with respect to the "y" -axis.

Points "(-\dfrac{\pi}{2}+2\pi n,\dfrac{\pi}{2}+2\pi n),n\in \Z" lie on the graph of "f(x)=-x"


Points "(\dfrac{\pi}{2}+2\pi n,\dfrac{\pi}{2}+2\pi n),n\in \Z" lie on the graph of "f(x)=x"


Points "(-\dfrac{3\pi}{2}+2\pi n,-\dfrac{3\pi}{2}+2\pi n),n\in \Z" lie on the graph of "f(x)=x"


Points "(\dfrac{3\pi}{2}+2\pi n,-\dfrac{3\pi}{2}+2\pi n),n\in \Z" lie on the graph of "f(x)=-x"



(ii)

The function "f(x)=\dfrac{\sin x}{x}" is not defined at "x=0."



"\lim\limits_{x\to 0}f(x)=\lim\limits_{x\to 0}\dfrac{\sin x}{x}=1"

The function "f(x)=\dfrac{\sin x}{x}" has a removable discontinuity at "x=0."

The function "f(x)=\dfrac{\sin x}{x}" is even. The graph is symmetric with respect to the "y" -axis.

"y\to0" as "x\to\pm \infin."


The graph of "f(x)=\dfrac{\sin x}{x}" is decaying oscillations (oscillations of continuously decreasing amplitude). The oscillations never stop, but go on decreasing in strength. The amplitude of oscillations at any point "x\not=0" is "1\/x."


3.



From right triangle


"h=L\sin \theta""h=23\sin 80\degree \ ft\approx22.65\ ft"